# Dynamic Rank Adjustment in Diffusion Policies for Efficient and Flexible Training

### v2

@@ Abstract
Diffusion policies trained via offline behavioral cloning have recently gained traction in robotic motion generation. While effective, these policies typically require a large number of trainable parameters. This model size affords powerful representations but also incurs high computational cost during training. Ideally, it would be beneficial to dynamically adjust the trainable portion as needed, balancing representational power with computational efficiency. For example, while overparameterization enables diffusion policies to capture complex robotic behaviors via offline behavioral cloning, the increased computational demand makes online interactive imitation learning impractical due to longer training time. To address this challenge, we present a framework, called DRIFT, that uses the Singular Value Decomposition to enable dynamic rank adjustment during diffusion policy training. We implement and demonstrate the benefits of this framework in DRIFT-DAgger, an imitation learning algorithm that can seamlessly slide between an offline bootstrapping phase and an online interactive phase. We perform extensive experiments to better understand the proposed framework, and demonstrate that DRIFT-DAgger achieves improved sample efficiency and faster training with minimal impact on model performance.
@@ I Introduction
  • We provide open-source implementations in Pytorch for the DRIFT framework and DRIFT-DAgger algorithm. [^1] We are in the progress of preparing the code for release.
@@ II-B Ranks in Diffusion Models
In the context of a diffusion policy that employs a U-Net with one-dimensional convolutional blocks as its network backbone, a weight matrix for each convolutional block $W\in\mathbb{R}^{n\times m}$ can be created by reshaping a corresponding weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times k}$ via
$W=\texttt{reshape}(W_{\text{conv}},(n,m)),$
where $C_{\text{out}}$ is the number of output channels, $C_{\text{in}}$ is the number of input channels, and $k$ is the kernel size. The reshaping can be performed by setting $n=C_{\text{out}}*k$ and $m=C_{\text{in}}$ or other equivalent view transformations.
@@ III-A Overparameterization and Intrinsic Rank
To tackle these challenges, researchers have observed that overparameterized models often reside in a low-dimensional subspace [1, 22]. This insight has led to techniques like Low-Rank Adaptation (LoRA) [12, 5], which fine-tune a small low-rank adapter while keeping the main model frozen. Although LoRA effectively reduces computational costs, it is primarily suited for fine-tuning pre-trained models and is less practical for training models from scratch due to its fixed low-rank structure [25].
@@ IV-B Rank Modulation
To address these limitations, we propose rank modulation, which leverages SVD to decompose weight matrices into components designated as either frozen or trainable ranks. More specifically, consider a weight matrix $W\in\mathbb{R}^{n\times m}$ that can be created from a corresponding weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times k}$ for a one-dimensional convolutional layer by reshaping, e.g., $C_{\text{out}}*k$ becomes $n$ and $C_{\text{in}}$ becomes $m$ (as covered in §II-B). Using this matrix, we first apply the SVD:
@@
where $U\in\mathbb{R}^{n\times n},\,\Sigma\in\mathbb{R}^{n\times m},\,V\in\mathbb{R}^{m\times m}.$ $U$ and $V$ are orthonormal matrices that represent rotations or reflections, while $\Sigma$ is a diagonal matrix containing scaling factors along the principal components of the weight matrix. Next, we split $U,\Sigma,$ and $V$ at a specified rank $r$ to partition trainable and frozen part of each matrix:
@@
$\Sigma$ $=\begin{bmatrix}\Sigma_{\text{train}}&0_{r\times(m-r)}\\ 0_{(n-r)\times r}&\Sigma_{\text{frozen}}\end{bmatrix},$
@@
$W_{\text{frozen}}$ $=U_{\text{frozen}}\,\Sigma_{\text{frozen}}\,V_{\text{frozen}}^{T}.$
During training, $\{U_{\text{train}},\Sigma_{\text{train}},V_{\text{train}}\}$ are the only parameters that receive gradient updates (rank-$r$ subspace), while $\{U_{\text{frozen}},\Sigma_{\text{frozen}},V_{\text{frozen}}\}$ remain fixed. Unlike LoRA, rank modulation performs a single convolution using the full $W=W_{\text{train}}+W_{\text{frozen}}$ via another simple view transformation in memory:
@@
Hence, the forward time complexity remains the same as a standard convolution. Because no additional new parameter is introduced during the training process, rank modulation can also preserve stable update even when the number of trainable ranks changes.
@@ IV-C Rank Scheduler
The rank scheduler uses a decay function to compute the current number of trainable ranks $r_{i}$ based on the current training epoch index $i$, and the maximum and minimum trainable ranks, $r_{\text{max}}$ and $r_{\text{min}}$. Once $r_{i}$ is determined, the trainable ranks are adjusted depending on the low-rank adapters. For instance, with LoRA, this process involves merging the current LoRA blocks and reinstantiating new blocks with the updated trainable ranks. In the case of rank modulation, this process repartitions $W_{\text{train}}$ and $W_{\text{frozen}}$.
@@ 7             $\pi_{N_{i}}=\texttt{Rank Reduction}(r_{i},\pi_{N_{i}})$
9if not use rank scheduler then
@@ VI-A Decay Functions
To assess performance and training efficiency when using different decay functions, we employ the success rate of a policy to evaluate the overall performance, where a high success rate indicates better completion of a given task. We also monitor the step loss during training as an indicator of convergence. Additionally, we measure the mean batch training time for each epoch, which includes the offline stage, online stage, and both stages combined, to reflect training efficiency when considered alongside the success rate and step loss.
@@ VI-B Terminal Rank
Figure 5: Summary of experimental results on mean batch training time (MBT) and success rate with different values for terminal ranks for DRIFT-DAgger.
@@
As shown in Fig. 5 and summarized in Table 5, decreasing the terminal ranks reduces training time but also impacts policy performance, as reflected in the success rate. This performance degradation occurs due to the diminished representational power of the model when the number of trainable ranks is reduced. Compared to reduced-rank methods for fine-tuning, such as [12], where fine-tuning with an adapter requires only 4 trainable ranks, DRIFT-DAgger with reduced-rank training from scratch necessitates a significantly higher number of trainable ranks, given that there is a noticeable performance drop when $r_{\text{min}}$ is set to 64. This is to fully leverage the benefits of both better representation power from overparameterization and improved computational efficiency from intrinsic low ranks, as fine-tuning with extremely small adapters does not capture the full representational potential of the model.
Conversely, setting $r_{\text{min}}$ too high, such as 512, offers no clear benefit in terms of policy performance, as the model already achieves a perfect success rate. The approximation error, as indicated by the loss curve in Fig. 5, also shows negligible differences between $r_{\text{min}}$ values of 256 and 512.
TABLE II: The configurations of DRIFT-DAgger for all simulation and real-world tasks. For simulation, we use other neural network policies trained with BC as experts, and compare the cosine similarity of the expert action and learner action with the threshold to determine whether the expert take over or not.
@@
Real-World Tasks Real-BS 150 150 50 -
@@ VI-C Benchmark Comparison
TABLE III: Summary of experimental results from simulation scenarios. The metrics include success rate (SR), mean and standard deviation of task duration (MSD), number of expert labels (NEL), and cumulative training time (CT). CT is measured in hours, MSD is measured in steps and at the scale of ×$\times 10^{2}$, and NEL is at the scale of ×$\times 10^{4}$
@@
The methods we evaluate in this benchmark comparison include Behavior Cloning (BC), HG-DAgger, and three variants of DRIFT-DAgger: one that uses LoRA, one that uses LoRA with rank scheduler, and one that uses rank modulation and rank scheduler. For methods utilizing rank scheduler, we apply the sigmoid decay function with steepness $\tau$ set to 0.5. We set $r_{\text{max}}$ and $r_{\text{min}}$ to 2048 and 256, respectively, based on the maximum rank of the diffusion policy and prior ablation study on the terminal rank. For all methods that use LoRA, we set the scaling factor $\alpha$ to 1.0. Experimental parameters related to the interactive mechanism, including the number of offline rollouts in $\mathcal{D}_{B}$, bootstrapping epochs $\mathcal{I}$, and online iterations $\mathcal{J}$, are provided in Table II.
@@
To fairly and efficiently evaluate different methods, for each scenario, we first train an expert policy using human-collected data from Mandlekar et al. [27] and Sun et al. [40] for robosuite and MVS tasks, respectively. The expert policy performs interventions when the cosine similarity between the learner actions and expert actions falls below a threshold, as detailed in Table II. The thresholds are computed based on the mean cosine similarity between consecutive steps in the expert training datasets.
As shown in Fig. 6 and Table III, DRIFT-DAgger variants demonstrate a pronounced reduction in cumulative training time compared to BC and HG-DAgger. Additionally, all interactive IL methods exhibit superior expert sample efficiency compared to BC, as evidenced by higher success rates with respect to the number of expert labels. An exception is observed in the DRIFT-DAgger variant with LoRA and rank scheduler, where the merging and re-injection of LoRA adapters destabilize training due to the initialization of new trainable parameters. In contrast, the DRIFT-DAgger variant that instantiates LoRA adapters only once during the transition to the online phase, or the variant that combines rank modulation and rank scheduler, achieve performance and sample efficiency comparable to HG-DAgger, which consistently uses full ranks instead of reduced ranks. The benefits of interactive IL are more pronounced in tasks with longer durations. Furthermore, the policies learned with DRIFT-DAgger using the combination of rank modulation and rank scheduler, exhibit good stability and better convergence to the desired behavior of the task, as indicated by the relatively lower standard deviation in task duration.
@@ VI-D Batch Training Time
TABLE IV: Summary of experimental results on mean batch training time (MBT) and success rate with full-rank and reduced-rank training methods.
@@
To better understand and demonstrate the improved training efficiency of DRIFT-DAgger, we conduct an ablation study on the mean batch training time across all stages, comparing different methods and variants. The methods evaluated in this ablation study include HG-DAgger and three DRIFT-DAgger variants: one using LoRA, one combining LoRA with a rank scheduler, and one employing rank modulation alongside rank scheduler. We use HG-DAgger as the baseline for full-rank training.
@@
The results are presented in Fig. 7 and Table IV. We observe that the DRIFT-DAgger variant with fixed-rank LoRA and the variant with rank modulation and rank scheduler both achieve success rates comparable to the full-rank baseline, HG-DAgger. When considering batch training time, the variant with rank modulation and rank scheduler reduces the mean batch training time across all stages by 11%, demonstrating improved training efficiency without sacrificing performance. Specifically, during the online training stage, this variant achieves an 18% reduction in batch training time compared to the full-rank baseline.
@@ VII Real-World Evaluation
TABLE V: Summary of experimental results from real-world scenarios. The metrics include success rate (SR), mean and standard deviation of task duration (MSD), number of expert labels (NEL), and cumulative training time (CT). CT is measured in hours, MSD measured in minutes, and NEL is at the scale of ×$\times 10^{4}$.
@@
The real-world tasks, depicted in Fig. 9, include block stacking, which involves placing a non-cuboid block on top of another; drawer assembling, which involves inserting two drawer boxes into a drawer container; and drawer interaction, which requires the agent to first get rid of a visual occlusion (cardboard box), grasp a red cube from a cluttered area, place it in the drawer, and then close the drawer. The blocks and drawer components are 3D-printed using models from [20] and [8]. Training configurations for each task are detailed in Table II, and the trained policies from each method are evaluated over 30 rollouts.
@@
Results, summarized in Table V, show trends consistent with the simulation experiments. All DRIFT-DAgger variants show noticeable reduction in cumulative training time. Interactive IL methods, except the DRIFT-DAgger variant with LoRA and rank scheduler, achieve similar or superior performance to BC, with improved sample efficiency, as indicated by a reduced number of expert labels. Despite using reduced-rank training, the other two DRIFT-DAgger variants perform comparably to HG-DAgger, which trains in a full-rank manner. The advantage of interactive methods becomes more pronounced for longer-duration tasks. For example, in the block stacking task, which has a mean duration of around 0.45 minutes, interactive methods improve sample efficiency by 11.32% to 13.17%. In contrast, for the drawer assembling task, with a mean duration of approximately 1.72 minutes, sample efficiency improves by 14.82% to 17.14%.

### v3

@@ Abstract
Diffusion policies trained via offline behavioral cloning have recently gained traction in robotic motion generation. While effective, these policies typically require a large number of trainable parameters. This model size affords powerful representations but also incurs high computational cost during training. Ideally, it would be beneficial to dynamically adjust the trainable portion as needed, balancing representational power with computational efficiency. For example, while overparameterization enables diffusion policies to capture complex robotic behaviors via offline behavioral cloning, the increased computational demand makes online interactive imitation learning impractical due to longer training time. To address this challenge, we present a framework, called DRIFT, that uses the Singular Value Decomposition to enable dynamic rank adjustment during diffusion policy training. We implement and demonstrate the benefits of this framework in DRIFT-DAgger, an imitation learning algorithm that can seamlessly slide between an offline bootstrapping phase and an online interactive phase. We perform extensive experiments to better understand the proposed framework, and demonstrate that DRIFT-DAgger achieves improved sample efficiency and faster training with minimal impact on model performance. The project website is available at: https://apollo-lab-yale.github.io/25-RSS-DRIFT-website/.
@@ I Introduction
  • We provide open-source implementations in Pytorch for the DRIFT framework and DRIFT-DAgger algorithm. [^1] https://github.com/Apollo-Lab-Yale/drift_dagger
@@ II-B Ranks in Diffusion Models
In the context of a diffusion policy that employs a U-Net with one-dimensional convolutional blocks as its network backbone, a weight matrix for each convolutional block $W\in\mathbb{R}^{m\times n}$ can be created by reshaping a corresponding weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times k}$ via
$W=\texttt{reshape}(W_{\text{conv}},(m,n)),$
where $C_{\text{out}}$ is the number of output channels, $C_{\text{in}}$ is the number of input channels, and $k$ is the kernel size. The reshaping can be performed by setting $m=C_{\text{out}}*k$ and $n=C_{\text{in}}$ or other equivalent view transformations.
@@ III-A Overparameterization and Intrinsic Rank
To tackle these challenges, researchers have observed that overparameterized models often reside in a low-dimensional subspace [1, 24, 47]. Pre-deep learning approaches like dynamical low-rank approximation [20] assume a derivable target matrix, unsuitable for deep learning where the target weight matrix is unknown. More modern techniques like Low-Rank Adaptation (LoRA) [13, 6] fine-tune a small low-rank adapter while keeping the main model frozen. Although LoRA and its variants like DyLoRA [43] and QLoRA [6] effectively reduces computational costs, it is primarily suited for fine-tuning pre-trained models and is less practical for training models from scratch due to its fixed low-rank structure [27].
@@ IV-B Rank Modulation
To address these limitations, we propose rank modulation, which leverages an SVD structure to decompose weight matrices into components designated as either frozen or trainable ranks. More specifically, consider a weight matrix $W\in\mathbb{R}^{m\times n}$ that can be created from a corresponding weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times k}$ for a one-dimensional convolutional layer by reshaping, e.g., $C_{\text{out}}*k$ becomes $m$ and $C_{\text{in}}$ becomes $n$ (as covered in §II-B). This matrix can be refactored into three matrices using the SVD:
@@
where $U\in\mathbb{R}^{m\times m},\,\Sigma\in\mathbb{R}^{m\times n},\,V\in\mathbb{R}^{n\times n}.$ $U$ and $V$ are orthonormal matrices that represent rotations or reflections, while $\Sigma$ is a diagonal matrix containing scaling factors. We can further split $U,\Sigma,$ and $V$ at a specified rank $r$ to partition trainable and frozen part of each matrix:
@@
$\Sigma$ $=\begin{bmatrix}\Sigma_{\text{train}}&0_{r\times(n-r)}\\ 0_{(m-r)\times r}&\Sigma_{\text{frozen}}\end{bmatrix},$
@@
$W_{\text{frozen}}$ $=U_{\text{frozen}}\,\Sigma_{\text{frozen}}\,V_{\text{frozen}}^{T},$
where $\Sigma_{\text{frozen}}$ holds smaller singular values than $\Sigma_{\text{train}}$. During training, $\{U_{\text{train}},\Sigma_{\text{train}},V_{\text{train}}\}$ are the only parameters that receive gradient updates (rank-$r$ subspace), while $\{U_{\text{frozen}},\Sigma_{\text{frozen}},V_{\text{frozen}}\}$ remain fixed.
The procedure described above, converting $W_{\text{conv}}$ into $W_{\text{train}}$ and $W_{\text{frozen}}$ using SVD and partitioning, is performed at the start of each epoch if the number of trainable ranks has changed since the previous epoch. This procedure also reorthonormalizes the $U$ and $V^{\top}$ matrices. While it is possible to reorthonormalize the $U$ and $V^{\top}$ matrices more frequently (such as after each gradient update) using QR decomposition, preliminary experiments showed that this introduced significant computational overhead without measurable performance gains (see Appendix §IX-A). Therefore, our implementation defaults to reorthonormalizing only via full SVD at the start of an epoch whenever the number of trainable ranks changes.
Unlike LoRA, rank modulation performs a single convolution using the full $W=W_{\text{train}}+W_{\text{frozen}}$ via another simple view transformation in memory:
@@
Hence, the forward time complexity remains the same as a standard convolution. Because no additional new parameters are introduced during the training process, rank modulation can also preserve stable updates even when the number of trainable ranks changes.
@@ IV-C Rank Scheduler
The rank scheduler uses a decay function to compute the current number of trainable ranks $r_{i}$ based on the current training epoch index $i$, and the maximum and minimum trainable ranks, $r_{\text{max}}$ and $r_{\text{min}}$. Once $r_{i}$ is determined, the trainable ranks are adjusted depending on the low-rank adapters. For instance, with LoRA, this process involves merging the current LoRA blocks and reinstantiating new blocks with the updated trainable ranks. In the case of rank modulation, this process entails computing the SVD of $W$ to produce updated $W_{\text{train}}$ and $W_{\text{frozen}}$ matrices.
@@ 7             $\pi_{N_{i}}=\texttt{Rank Reduction}(r_{i},\pi_{N_{i}})$
9      if not use rank scheduler then
@@ VI-A Decay Functions
To assess performance and training efficiency across different decay functions, we evaluate the policy’s success rate, with higher values indicating better task completion. We also track the step loss during training as a measure of convergence. To assess training efficiency, we record the mean batch training time per epoch, separately for the offline stage, online stage, and their combination, interpreting it alongside the success rate and step loss.
@@ VI-B Terminal Rank
TABLE II: Summary of experimental results on mean batch training time (MBT) and success rate with different values for terminal ranks for DRIFT-DAgger.
@@
As shown in Fig. 4 and summarized in Table II, decreasing the terminal ranks reduces training time but also impacts policy performance, as reflected in the success rate. This performance degradation occurs due to the diminished representational power of the model when the number of trainable ranks is reduced. Compared to reduced-rank methods for fine-tuning, such as [13], where fine-tuning with an adapter requires only 4 trainable ranks, DRIFT-DAgger with reduced-rank training from scratch necessitates a significantly higher number of trainable ranks, given that there is a noticeable performance drop when $r_{\text{min}}$ is set to 64. This is to fully leverage the benefits of both better representation power from overparameterization and improved computational efficiency from intrinsic low ranks, as fine-tuning with extremely small adapters does not capture the full representational potential of the model.
Conversely, setting $r_{\text{min}}$ too high, such as 512, offers no clear benefit in terms of policy performance, as the model already achieves a perfect success rate. The approximation error, as indicated by the loss curve in Fig. 4, also shows negligible differences between $r_{\text{min}}$ values of 256 and 512. We also run $r_{\text{min}}$ sweep for another MVS task, and the results follow similar trends as the pick-and-place scenario that $r_{\text{min}}=256$ provides the best balance of efficiency and performance (see Appendix §IX-B).
TABLE III: The configurations of DRIFT-DAgger for all simulation and real-world tasks. For simulation, we use other neural network policies trained with BC as experts, and compare the cosine similarity of the expert action and learner action with the threshold to determine whether the expert take over or not.
@@
Real Real-BS 150 150 50 -
@@ VI-C Benchmark Comparison
TABLE IV: Summary of experimental results from simulation scenarios. The metrics include success rate (SR), mean and standard deviation of task duration (MSD), number of expert labels (NEL), and cumulative training time (CT). CT is measured in hours, MSD is measured in steps and at the scale of ×$\times 10^{2}$, and NEL is at the scale of ×$\times 10^{4}$
@@
The methods we evaluate in this benchmark comparison include Behavior Cloning (BC), HG-DAgger, and three variants of DRIFT-DAgger: one that uses LoRA, one that uses LoRA with rank scheduler, and one that uses rank modulation and rank scheduler. For methods utilizing rank scheduler, we apply the sigmoid decay function with steepness $\tau$ set to 0.5. We set $r_{\text{max}}$ and $r_{\text{min}}$ to 2048 and 256, respectively, based on the maximum rank of the diffusion policy and prior ablation study on the terminal rank. For all methods that use LoRA, we set the scaling factor $\alpha$ to 1.0. Experimental parameters related to the interactive mechanism, including the number of offline rollouts in $\mathcal{D}_{B}$, bootstrapping epochs $\mathcal{I}$, and online iterations $\mathcal{J}$, are provided in Table III.
@@
To fairly and efficiently evaluate different methods, for each scenario, we first train an expert policy using human-collected data from Mandlekar et al. [29] and Sun et al. [42] for robosuite and MVS tasks, respectively. The expert policy performs interventions when the cosine similarity between the learner actions and expert actions falls below a threshold, as detailed in Table III. The thresholds are computed based on the mean cosine similarity between consecutive steps in the expert training datasets.
As shown in Fig. 5 and Table IV, DRIFT-DAgger variants demonstrate a pronounced reduction in cumulative training time compared to BC and HG-DAgger. Additionally, all interactive IL methods exhibit superior expert sample efficiency compared to BC, as evidenced by higher success rates with respect to the number of expert labels. An exception is observed in the DRIFT-DAgger variant with LoRA and rank scheduler, where the merging and re-injection of LoRA adapters destabilize training due to the initialization of new trainable parameters. In contrast, the DRIFT-DAgger variant that instantiates LoRA adapters only once during the transition to the online phase, or the variant that combines rank modulation and rank scheduler, achieve performance and sample efficiency comparable to HG-DAgger, which consistently uses full ranks instead of reduced ranks. The benefits of interactive IL are more pronounced in tasks with longer durations. Furthermore, the policies learned with DRIFT-DAgger using the combination of rank modulation and rank scheduler, exhibit good stability and better convergence to the desired behavior of the task, as indicated by the relatively lower standard deviation in task duration.
@@ VI-D Batch Training Time
TABLE V: Summary of experimental results on mean batch training time (MBT) and success rate with full-rank and reduced-rank training methods.
@@
To better understand and demonstrate the improved training efficiency of DRIFT-DAgger, we conduct an ablation study on the mean batch training time across all stages. The methods evaluated in this ablation study include HG-DAgger and three DRIFT-DAgger variants: one using LoRA, one combining LoRA with a rank scheduler, and one employing rank modulation alongside rank scheduler. We use HG-DAgger as the baseline for full-rank training.
@@
The results are presented in Fig. 6 and Table V. We observe that the DRIFT-DAgger variant with fixed-rank LoRA and the variant with rank modulation and rank scheduler both achieve success rates comparable to the full-rank baseline, HG-DAgger. When considering batch training time, the variant with rank modulation and rank scheduler reduces the mean batch training time across all stages by 11%, demonstrating improved training efficiency without sacrificing performance. Specifically, during the online training stage, this variant achieves an 18% reduction in batch training time compared to the full-rank baseline.
@@ VII Real-World Evaluation
TABLE VI: Summary of experimental results from real-world scenarios. The metrics include success rate (SR), mean and standard deviation of task duration (MSD), number of expert labels (NEL), and cumulative training time (CT). CT is measured in hours, MSD measured in minutes, and NEL is at the scale of ×$\times 10^{4}$.
@@
The real-world tasks, depicted in Fig. 8, include block stacking, which involves placing a non-cuboid block on top of another; drawer assembling, which involves inserting two drawer boxes into a drawer container; and drawer interaction, which requires the agent to first get rid of a visual occlusion (cardboard box), grasp a red cube from a cluttered area, place it in the drawer, and then close the drawer. The blocks and drawer components are 3D-printed using models from [22] and [9]. Training configurations for each task are detailed in Table III, and the trained policies from each method are evaluated over 30 rollouts.
@@
Results, summarized in Table VI, show trends consistent with the simulation experiments. All DRIFT-DAgger variants show noticeable reduction in cumulative training time. Interactive IL methods, except the DRIFT-DAgger variant with LoRA and rank scheduler, achieve similar or superior performance to BC, with improved sample efficiency, as indicated by a reduced number of expert labels. Despite using reduced-rank training, the other two DRIFT-DAgger variants perform comparably to HG-DAgger, which trains in a full-rank manner. The advantage of interactive methods becomes more pronounced for longer-duration tasks. For example, in the block stacking task, which has a mean duration of around 0.45 minutes, interactive methods improve sample efficiency by 11.32% to 13.17%. In contrast, for the drawer assembling task, with a mean duration of approximately 1.72 minutes, sample efficiency improves by 14.82% to 17.14%.